# Multi-mode Coherent Detection Ghost Imaging Lidar and Vibration-Mode Imaging

### v1

@@ 1 Introduction
Ghost imaging[1, 2, 3] has attracted significant interest in recent decades, as it can employ optical field fluctuation to encode and compress high-dimensional spatial image information into a lower-dimensional detection space that can be directly measurable by a detector, and allows the detection and imaging processes to be physically separated. Ghost imaging separates the tasks of optical signal acquisition and spatial resolution, enabling spatially resolved imaging performance comparable to that of detector arrays while operating within the spatiotemporal bandwidth product of a single detector. These advantages make it suitable for non-traditional imaging scenarios, including staring imaging radar[4, 5, 6, 7, 8, 9], ultra-low dose X-ray imaging[10], electron[11], and neutron radiographic imaging[12, 13], as well as atomic ghost imaging [14], snapshot multidimensional ghost imaging[15, 16], super-resolution microscopic based on spectral ghost imaging[17], and scattering imaging[18, 19, 20, 21, 22, 23], etc. Besides implementations based on intensity-correlations only can reconstruct the target’s intensity image, subsequent developments in field correlation ghost imaging employing homodyne detection have shown that both quantum-entangled light sources[24] and classical thermal light sources[25] are capable of retrieving the complex reflectivity of the target.
@@
With recent progress in optical coherent detection[26, 27, 28, 29], it can now achieve sensitivity close to the shot-noise limit, just like intensity detection, and can effectively suppress background light interference while acquiring amplitude and phase information of the light field. These capabilities have motivated the development of coherent detection imaging lidar systems, notably: (i) synthetic aperture imaging lidar (SAIL), which achieves high spatial resolution through synthetic aperture synthesis[30, 31, 32, 33, 34], and (ii) coherent focal-plane array (FPA) imaging lidar, which enables real-time parallel detection with high detection efficiency[35, 36, 37, 38, 39, 40, 41]. However, the fundamental limitation of SAIL is that it relies on platform-induced relative motion to modulate the return signal. Coherent FPA imaging, on the other hand, is challenged by the large spatiotemporal bandwidth product required for both detection and data processing[42, 43].
@@
The integration of ghost imaging and coherent detection provides a promising alternative to the two previously mentioned coherent-detection–based imaging lidar systems[44, 45], which can retrieve spatial resolution by field correlation with the detected signal from single pixel point detector, relax the spatiotemporal bandwidth requirements on the receiver, and enable staring imaging without platform-induced relative motion. In this work, through theoretical analysis and experiments, we show that in the staring imaging mode, introducing bucket-detector-based multi-mode CD-GI lidar system breaks the constraint imposed by Siegman antenna theorem via field correlation to decouple the coherent summation of reflected multi-mode light fields and enables the simultaneous retrieval of both amplitude and phase information of the target, thereby achieving complex reflectivity imaging. Moreover, it allows access to the spatial distribution of targets’ vibration modes in dynamic scenes within the field of view (FoV), thereby providing a approach for low spatiotemporal bandwidth product and staring imaging.
@@ 2.1 Light field transmission and coherent detection
The spatiotemporally modulated light field is generated by the coordinated temporal and spatial modulation modules illustrated in Fig. 1. A narrow-linewidth laser serves as the seed source, whose emitted light field can be written as $E_{0}=A_{0}\exp{(j2\pi f_{c}t)}$, with a center frequency of $f_{c}$. To achieve distance resolution, the seed light is first modulated by an electro-optic modulator (EOM), which is driven by a chirped radio-frequency signal $s(t)$. The modulated beam is subsequently split by a fiber beam splitter into an object arm and a LO arm. In the object arm, an acousto-optic modulator (AOM) introduces a frequency shift $f_{if}$. Subsequently, the temporally modulated beam is collimated and incident onto a digital micromirror device (DMD) operating in an externally triggered mode synchronized with the temporal modulation period. The DMD is intrinsically a binary amplitude modulator, and generates a series of predefined spatial encoding patterns noted as $E_{s}^{(i)}(\boldsymbol{\rho}_{s})=A_{s}^{(i)}\left(\boldsymbol{\rho}_{s}\right)\exp\left[j\phi_{s}^{(i)}\left(\boldsymbol{\rho}_{s}\right)\right]$. Since the DMD is a pre-programmed encoding device acting only on the object arm, the reference-arm field can be expressed as a virtual encoding: $E_{r}^{(i)}(\boldsymbol{\rho}_{r})=A_{s}^{(i)}\left(\boldsymbol{\rho}_{r}\right)\exp\left[j\phi_{s}^{(i)}\left(\boldsymbol{\rho}_{r}\right)\right]$. This process generates a spatiotemporally modulated transmitted field, which can be represented as:
@@
After propagation through a imaging transmitting system, as shown in Fig. 1, the modulated field illuminates the target, which is modeled by a spatially varying complex reflectivity combined with a time-dependent phase term accounting for range variations or surface vibrations:
@@
where $t_{o}(\boldsymbol{\rho}_{o})$ indicates the amplitude reflectivity of target, $\phi_{o}(\boldsymbol{\rho}_{o})$ represents the phase of target and $z(\boldsymbol{\rho}_{o},t)$ indicates the distance variation caused by the micro-vibration of target. The resulting multi-mode back-scattered field is collected by the imaging receiving system and coherently interfered with the LO in a spatial optical mixer, followed by the balanced bucket detector. The architecture of this multi-mode detection module is described in Fig. 1.
The reflected light fields on the surface of the detector can be represented as (see Supplement 1, Sec. S1 for detailed derivation process):
@@
where $k=2\pi/\lambda$ represents the wave number; $z_{1}$ and $z_{2}$ denote the object distance and image distance of the imaging transmitting system; $z_{3}$ and $z_{4}$ denote the object and image distances of the reception imaging system; $\exp[j(\phi_{i,1}+\phi_{i,2})]$ represents the overall phase distortion introduced by atmospheric turbulence during the propagation of the $i$-th spatial modulated pattern, $\phi_{i,1}$ and $\phi_{i,2}$ represent the turbulence-induced phase perturbations along the forward and return propagation paths; $D_{f}$ represents the aperture diameter of the receiving lens; $A_{o}$ represents the target area; $\boldsymbol{\rho}_{o}=(x_{o},y_{o})$ denotes the transverse coordinates on the target plane and $\boldsymbol{\rho}_{d}=(x_{d},y_{d})$ denotes the transverse coordinates on the detector plane. After balanced detection, the optical signal is converted into an electrical signal $i^{(i)}(t)$:
@@
Here, $\tau=(z_{1}+z_{2}+z_{3}+z_{4})/{c}$ is the time delay caused by the target distance; $R$ denotes the responsivity of the detector; $c.c.$ denotes the complex conjugate of $E_{d}^{(i)}(\boldsymbol{\rho}_{d},t)E_{\text{Lo-mod}}^{*}(t)$ and $A_{d}$ represents the area of the detector.
@@ 2.2 Signal processing and Vibration-Mode imaging reconstruction
A static target can be regarded as a target with vibration frequency of zero. To obtain an explicit expression for the IF signal, the Fast Fourier transform is applied to $i^{{(i)}}(t)$ ($L$ is the total length of the signal.), yielding (see Supplement 1, Sec. S2 for detailed derivation process):
$\widetilde{i^{(i)}}\left(f=f_{if}+\frac{B}{T}\,\tau\right)$ $=\frac{L}{2}A_{0}^{2}\pi R\frac{\sqrt{\eta_{\mathrm{Sig}}}\sqrt{\eta_{\mathrm{Lo}}}}{\lambda^{4}z_{1}z_{2}z_{3}z_{4}}\exp\left[jk(z_{1}+z_{2}+z_{3}+z_{4})+j2\pi\left(f_{0}\tau-\frac{B}{2T}\tau^{2}\right)\right]$ (6)
$\times\int_{A_{d}}\exp\left(\frac{j\pi}{\lambda z_{4}}|\boldsymbol{\rho}_{d}|^{2}\right)\int_{A_{o}}\exp\left[\frac{j\pi}{\lambda}\left(\frac{z_{1}+z_{2}}{z_{2}^{2}}+\frac{1}{z_{3}}\right)|\boldsymbol{\rho}_{o}|^{2}\right]$
$\times\frac{J_{1}\left[\frac{D_{f}\pi}{\lambda}\left(\frac{\boldsymbol{\rho}_{o}}{z_{3}}+\frac{\boldsymbol{\rho}_{d}}{z_{4}}\right)\right]}{\frac{\pi}{\lambda}\left(\frac{\boldsymbol{\rho}_{o}}{z_{3}}+\frac{\boldsymbol{\rho}_{d}}{z_{4}}\right)}E_{s}^{(i)}(-\frac{z_{1}}{z_{2}}\boldsymbol{\rho}_{o})\tilde{T}(\boldsymbol{\rho}_{o})d\boldsymbol{\rho}_{o}d\boldsymbol{\rho}_{d}\cdot\exp\left(j\phi_{i}\right),$
where $\phi_{i}=\phi_{i,1}+\phi_{i,2}$. Due to the phase disturbances arising during the detection process, an appropriate phase correction is required to recover the IF signal:
$\widetilde{i}_{\mathrm{correct}}^{(i)}\left(f=f_{if}+\frac{B}{T}\,\tau\right)=\frac{L}{2}A_{0}^{2}\pi R\frac{\sqrt{\eta_{\mathrm{Sig}}}\sqrt{\eta_{\mathrm{Lo}}}}{\lambda^{4}z_{1}z_{2}z_{3}z_{4}}\cdot\frac{\displaystyle\int_{A_{o}}W(\boldsymbol{\rho}_{o})\,E_{s}^{(i)}\!\left(-\frac{z_{1}}{z_{2}}\boldsymbol{\rho}_{o}\right)\,\tilde{T}(\boldsymbol{\rho}_{o})\,d\boldsymbol{\rho}_{o}}{\displaystyle\exp{\left[j\cdot\mathrm{arg}(\int_{A_{o}}W(\boldsymbol{\rho}_{o})\,\tilde{T}(\boldsymbol{\rho}_{o})\,d\boldsymbol{\rho}_{o})\right]}},$ (7)
$W(\boldsymbol{\rho}_{o})=\exp\!\left[\frac{j\pi}{\lambda}\left(\frac{z_{1}+z_{2}}{z_{2}^{2}}+\frac{1}{z_{3}}\right)|\boldsymbol{\rho}_{o}|^{2}\right]C(\boldsymbol{\rho}_{o}),$ (7.1)
$C(\boldsymbol{\rho}_{o})=\int_{A_{d}}\exp\!\left(\frac{j\pi}{\lambda z_{4}}|\boldsymbol{\rho}_{d}|^{2}\right)\frac{J_{1}\left[\frac{D_{f}\pi}{\lambda}\left(\frac{\boldsymbol{\rho}_{o}}{z_{3}}+\frac{\boldsymbol{\rho}_{d}}{z_{4}}\right)\right]}{\frac{\pi}{\lambda}\left(\frac{\boldsymbol{\rho}_{o}}{z_{3}}+\frac{\boldsymbol{\rho}_{d}}{z_{4}}\right)}d\boldsymbol{\rho}_{d},$ (7.2)
where $(i)\in\mathbb{N}$ and $\mathrm{arg}(\cdot)$ denotes extracting the phase of a complex signal. The correction procedure, the detailed derivation of the corrected IF signal and the field correlation image of the complex reflectivity for a static target are provided in Supplement 1, Sec. S2.A.
In practical remote-sensing scenarios, many targets exhibit micro-vibration motions, such as engine-induced vibrations[46], rotor oscillations[47], and structural flutter[48], which arise from continuous external excitation and manifest as steady-state periodic displacements. In this work, the target is assumed to undergo continuous vibration throughout the sampling interval.
@@
where $\Delta z(\boldsymbol{\rho}_{o})$ is the vibration amplitude, $f_{v}$ is the vibration frequency and $\varphi_{o}(\boldsymbol{\rho}_{o})$ is the initial vibration phase.
Since the experimental system can extract the target’s time frequency information from the Short-time Fourier transform (STFT) spectrum[50, 51, 52, 53]. To ensure that the spatiotemporal modulation on the DMD is synchronized with the target vibration, the repetition rate of the spatial modulation sequence on the DMD is set to be an integer divisor of the vibration frequency. Under this condition, assuming the total length of the signal $i^{(i)}(t)$ is $L$, a sliding rectangular window function $w(t)$ is used with length $l$ and overlap $d$, $k$ is the number of the corresponding windows: $k=0,1,2,\dots,\left\lfloor\frac{L-l}{l-d}\right\rfloor$. After performing field correlation of the corrected complex signal $\widetilde{i}_{\mathrm{correct}-k}^{(i)}\!\left(f=\omega_{k}^{(h)}/2\pi\right)$, which corresponds to the $h$-th micro-Doppler frequency in the $k$-th time slice of the STFT spectrum, with the reference-arm field $\left[E_{r}^{(i)}(\boldsymbol{\rho}_{r})\right]^{*}$, we can obtain the spatial distribution corresponding to the micro-Doppler frequency for $f=\omega_{k}^{(h)}/2\pi$ of the target:
@@
Here, the ensemble averaging $\langle\cdot\rangle$ is performed over the encoded optical field $i$ corresponding to the $h$-th micro-Doppler frequency component within the $k$-th time window. The vibration amplitude corresponding to the target spatial position can be converted as follows (see Supplement 1, Sec. S2.B for detailed derivation process):
@@
Field correlation reconstruction is sequentially performed over the $k$ time windows defined by the STFT. The spatial images reconstructed from these windows are combined with the corresponding vibration amplitude to recover the target’s vibration modes.

### v2

@@ 1 Introduction
Ghost imaging[1, 2, 3] has attracted significant interest in recent decades, as it can employ optical field fluctuation to encode and compress high-dimensional spatial image information into a lower-dimensional detection space that can be directly measurable by a detector, and allows the detection and imaging processes to be physically separated. Ghost imaging separates the tasks of optical signal acquisition and spatial resolution, enabling spatially resolved imaging performance comparable to that of detector arrays while operating within the spatiotemporal bandwidth product of a single detector. These advantages make it suitable for non-traditional imaging scenarios, including staring imaging radar and lidar[4, 5, 6, 7, 8, 9], ultra-low dose X-ray imaging[10], electron[11], and neutron radiographic imaging[12, 13], as well as atomic ghost imaging [14], snapshot multidimensional ghost imaging[15, 16], super-resolution imaging based on spectral dimensional information[17], and scattering imaging[18, 19, 20, 21, 22, 23], etc. While implementations based on intensity-correlations only can reconstruct the target’s intensity image, subsequent developments in field correlation ghost imaging employing homodyne detection have shown that both quantum-entangled light sources[24] and classical thermal light sources[25] are capable of retrieving the complex reflectivity of the target.
@@
Recent advancements in optical coherent detection[26, 27, 28, 29] have enabled sensitivities approaching the shot-noise limit, comparable to intensity detection, while simultaneously offering the capability to suppress background light interference and retrieve both the amplitude and phase information of the light field. These capabilities have motivated the development of coherent detection imaging lidar systems, notably: (i) synthetic aperture imaging lidar (SAIL), which circumvents the diffraction limit to achieve high spatial resolution through synthetic aperture synthesis[30, 31, 32, 33, 34], and (ii) coherent focal-plane array (FPA) imaging lidar, which enables real-time parallel detection with high detection efficiency[35, 36, 37, 38, 39, 40, 41]. However, the fundamental limitation of SAIL is that it relies on platform-induced relative motion to modulate the return signal. Concurrently, coherent FPA imaging is challenged by the spatiotemporal bandwidth product required for both large-scale detection and high-speed data processing[42, 43].
@@
The integration of ghost imaging and coherent detection provides a promising alternative to the aforementioned coherent-detection–based imaging lidar systems[44, 45], which can retrieve spatial resolution via field correlation using only a single-bucket detector, alleviate the spatiotemporal bandwidth requirements on the receiver, and enable staring imaging without platform-induced relative motion. In this work, through theoretical analysis and experiments, we show that the bucket-detector-based multi-mode CD-GI lidar system overcomes the limitation imposed by the Siegman antenna theorem through field correlation, decoupling the coherent summation of reflected multi-mode light fields. It enables the simultaneous retrieval of both amplitude and phase information of the target, and provides access to the spatial distribution of targets’ vibration modes in dynamic scenes within the field of view (FoV). Ultimately, our approach offers a framework for staring imaging with reduced spatiotemporal bandwidth product requirements.
@@ 2.1 Light field transmission and coherent detection
The spatiotemporally modulated light field is generated by the coordinated temporal and spatial modulation modules illustrated in Fig. 1. A narrow-linewidth laser serves as the seed source, yielding an emitted light field described by $E_{0}=A_{0}\exp{(j2\pi f_{c}t)}$, where $f_{c}$ denotes the center frequency. To achieve distance resolution capability, the seed light is initially modulated by an electro-optic modulator (EOM) driven by a chirped radio-frequency signal $s(t)$. The resultant beam is partitioned by a fiber beam splitter into an object arm and a LO arm. In the object arm, an acousto-optic modulator (AOM) induces a frequency shift $f_{if}$. Subsequently, the temporally modulated beam is collimated and incident onto a digital micromirror device (DMD) operating in an externally triggered mode, synchronized with the temporal modulation period. The DMD is intrinsically a binary amplitude modulator, and generates a series of predefined spatial encoding patterns denoted as $E_{s}^{(i)}(\boldsymbol{\rho}_{s})=A_{s}^{(i)}\left(\boldsymbol{\rho}_{s}\right)\exp\left[j\phi_{s}^{(i)}\left(\boldsymbol{\rho}_{s}\right)\right]$. Since the DMD is a pre-programmed encoding device acting only on the object arm, the reference-arm field can be defined through a virtual encoding formulated as: $E_{r}^{(i)}(\boldsymbol{\rho}_{r})=A_{s}^{(i)}\left(\boldsymbol{\rho}_{r}\right)\exp\left[j\phi_{s}^{(i)}\left(\boldsymbol{\rho}_{r}\right)\right]$. This architecture produces a spatiotemporally modulated transmitted field, which can be expressed as:
@@
After propagation through the imaging transmission module, as shown in Fig. 1, the modulated field illuminates the target, which is modeled by a spatially varying complex reflectivity combined with a time-dependent phase term accounting for surface vibrations:
@@
where $t_{o}(\boldsymbol{\rho}_{o})$ indicates the amplitude reflectivity of target, $\phi_{o}(\boldsymbol{\rho}_{o})$ represents the phase of target and $z(\boldsymbol{\rho}_{o},t)$ indicates the distance variation caused by the micro-vibration of target. The resulting multi-mode reflected light field is collected by the imaging receiving module and coherently interfered with the LO in a spatial optical mixer, followed by the balanced bucket detector. The architecture of this multi-mode detection module is described in Fig. 1. The reflected light fields on the surface of the detector can be represented as (see Supplement 1, Sec. S1 for detailed derivation process):
@@
where $k=2\pi/\lambda$ designates the wave number; $z_{1}$ and $z_{2}$ represent the object and image distances of the imaging transmission module, while $z_{3}$ and $z_{4}$ refer to the corresponding distances for the imaging receiving module. The term $\exp[j(\phi_{i,1}+\phi_{i,2})]$ accounts for the overall phase distortion introduced by atmospheric turbulence during the propagation of the $i$-th spatially modulated light field, with $\phi_{i,1}$ and $\phi_{i,2}$ characterizing the forward and return paths, respectively. Additionally, $D_{f}$ denotes the aperture diameter of the receiving lens, and $A_{o}$ signifies the target area. Finally, $\boldsymbol{\rho}_{o}=(x_{o},y_{o})$ stands for the transverse coordinates on the target plane, whereas $\boldsymbol{\rho}_{d}=(x_{d},y_{d})$ indicates the coordinates on the detector plane. After balanced detection, the optical signal is converted into an electrical signal $i^{(i)}(t)$:
@@
Here, $\tau=(z_{1}+z_{2}+z_{3}+z_{4})/{c}$ is the time delay due to the propagation of light; $R$ denotes the responsivity of the detector; $c.c.$ denotes the complex conjugate of $E_{d}^{(i)}(\boldsymbol{\rho}_{d},t)E_{\text{Lo-mod}}^{*}(t)$ and $A_{d}$ represents the area of the detector.
@@ 2.2 Signal processing and Vibration-Mode imaging reconstruction
In practical remote-sensing scenarios, many targets exhibit micro-vibration motions, such as engine-induced vibrations[46], rotor oscillations[47], and structural flutter[48], which arise from continuous external excitation.
@@
where $\Delta z(\boldsymbol{\rho}_{o})$ denotes the vibration height, $f_{v}$ is the vibration frequency and $\varphi_{o}(\boldsymbol{\rho}_{o})$ represents the initial vibration phase.
To simplify the problem, the drive frequency of target is set as an integer multiple of the DMD repetition rate, ensuring synchronization with the target vibration. The detected signal $i^{(i)}(t)$ is then subjected to a Short-Time Fourier Transform (STFT). The $i^{(i)}(t)$ is then subjected to a short-time Fourier transform (STFT). Assuming the total length of $i^{(i)}(t)$ is $L$, a sliding rectangular window function $w(t)$ is used with length $l$ and overlap $d$, $k$ is the number of the corresponding windows: $k=0,1,2,\dots,\left\lfloor\frac{L-l}{l-d}\right\rfloor$, the result of the STFT can be expressed as:
$\widetilde{i}_{k}^{(i)}\left(f=\frac{\omega_{k}^{(h)}}{2\pi}\right)$ $=Rl\int_{A_{d}}\int_{A_{o}}A_{1}(\boldsymbol{\rho}_{o},\boldsymbol{\rho}_{d})\pi e^{j\Phi(\boldsymbol{\rho}_{o},\boldsymbol{\rho}_{d})}e^{-j\omega_{k}^{(h)}\left[k(l-d)+\frac{l}{2}\right]}\mathrm{sinc}(\frac{\omega_{k}^{(h)}l}{2})d\boldsymbol{\rho}_{o}d\boldsymbol{\rho}_{d}$ (7)
$+Rl\int_{A_{d}}\int_{A_{o}}A_{1}(\boldsymbol{\rho}_{o},\boldsymbol{\rho}_{d})\pi e^{-j\Phi(\boldsymbol{\rho}_{o},\boldsymbol{\rho}_{d})}e^{j\omega_{k}^{(h)}\left[k(l-d)+\frac{l}{2}\right]}\mathrm{sinc}(\frac{\omega_{k}^{(h)}l}{2})d\boldsymbol{\rho}_{o}d\boldsymbol{\rho}_{d},$
$\Phi(\boldsymbol{\rho}_{o},\boldsymbol{\rho}_{d})$ $=k(z_{1}+z_{2}+z_{3}+z_{4})+\phi_{o}(\boldsymbol{\rho}_{o})+\phi_{s}^{(i)}\left(-\frac{z_{1}}{z_{2}}\boldsymbol{\rho}_{o}\right)+\frac{\pi}{\lambda z_{4}}|\boldsymbol{\rho}_{d}|^{2}$
$+\frac{\pi}{\lambda}\left(\frac{z_{1}+z_{2}}{z_{2}^{2}}+\frac{1}{z_{3}}\right)|\boldsymbol{\rho}_{o}|^{2}+2\pi\left(f_{0}\tau-\frac{B}{2T}\tau^{2}\right)+2kz_{0}(\boldsymbol{\rho}_{o})+\phi_{i}$ (7.2)
where $\omega_{k}^{(h)}=2\pi\left[f_{if}+\frac{B}{T}\tau+\frac{4\pi f_{v}\Delta z(\boldsymbol{\rho}_{o})}{\lambda}\cos\left[2\pi f_{v}\cdot k(l-d)+\phi_{o}(\boldsymbol{\rho}_{o})\right]\right]$ is the corresponding to the $h$-th micro-Doppler frequency in the $k$-th time slice of STFT.
Atmospheric turbulence correction has been a critical focus in coherent detection techniques, with many previous works [50, 51, 52, 53] addressing this issue. In this work, we also consider turbulence correction using a time-division strategy[54], where even frames correct the odd frames in the encoding process. The detailed derivation is provided in Supplement 1, Sec. S2.A. The corrected IF signal expression is given as follows:
$\widetilde{i}_{\mathrm{correct}-k}^{(i)}\left(f=\frac{\omega_{k}^{(h)}}{2\pi}\right)=\frac{\widetilde{i}_{k}^{(2i-1)}\left(f=\frac{\omega_{k}^{(h)}}{2\pi}\right)}{\exp{\left[j\cdot\mathrm{arg}(\widetilde{i}_{k}^{(2i)}\left(f=\frac{\omega_{k}^{(h)}}{2\pi}\right))\right]}}$ (8)
After performing field correlation of the corrected IF signal $\widetilde{i}_{\mathrm{correct}-k}^{(i)}\!\left(f=\omega_{k}^{(h)}/2\pi\right)$ with the reference-arm field $\left[E_{r}^{(i)}(\boldsymbol{\rho}_{r})\right]^{*}$, we can obtain the spatial distribution corresponding to the micro-Doppler frequency $f=\omega_{k}^{(h)}/2\pi$ of the target:
@@
Here, the ensemble averaging $\langle\cdot\rangle$ is performed over the encoded optical field $i$ corresponding to the $h$-th micro-Doppler frequency component within the $k$-th time window. The vibration height corresponding to the target spatial position can be converted as follows (see Supplement 1, Sec. S2.B for detailed derivation process):
@@
The spatial images $G_{k}^{(h)}(\boldsymbol{\rho}_{r})$ reconstructed from these windows are combined with the corresponding vibration height $A_{k}^{(h)}(\boldsymbol{\rho}_{r})$ to recover the target’s vibration modes.
A static target represents a special case where the vibration frequency vanishes ($f_{v}=0$), yielding zero micro-Doppler shifts across all temporal slices. In this scenario, the STFT processing degenerates into a standard FFT, and the complex reflectivity reconstruction result in Eq. (9) can be simplified as:
$G(\boldsymbol{\rho}_{r})$ $=\left\langle\widetilde{i}_{\mathrm{correct}}^{(i)}\left(f=f_{if}+\frac{B}{T}\,\tau\right)\cdot\left[E_{r}^{(i)}(\boldsymbol{\rho}_{r})\right]^{*}\right\rangle$ (11)
$\propto A_{0}^{2}\pi lR\frac{\sqrt{\eta_{\mathrm{Sig}}}\sqrt{\eta_{\mathrm{Lo}}}}{\lambda^{4}z_{1}z_{2}z_{3}z_{4}}\exp\left[jk(z_{1}+z_{2}+z_{3}+z_{4})\right]\exp\left[j2\pi\left(f_{0}\tau-\frac{B}{2T}\tau^{2}\right)\right]$
$\times\exp\left[\frac{j\pi}{\lambda}\left(\frac{z_{1}+z_{2}}{z_{2}^{2}}+\frac{1}{z_{3}}\right)|-\frac{z_{2}}{z_{1}}\boldsymbol{\rho}_{r}|^{2}\right]\tilde{T}(-\frac{z_{2}}{z_{1}}\boldsymbol{\rho}_{r})\frac{J_{1}\left[\frac{D_{f}\pi}{\lambda}\left(\frac{z_{2}}{z_{1}z_{3}}\boldsymbol{\rho}_{r}\right)\right]}{\frac{\pi}{\lambda}\left(\frac{z_{2}}{z_{1}z_{3}}\boldsymbol{\rho}_{r}\right)}.$
The derivation of the field correlation reconstruction for static targets is provided in Supplement 1, Sec. S2.A.